# Efficient Alignment of Unconditioned Action Prior for Language-conditioned Pick and Place in Clutter

### v2

@@ Abstract
We study the task of language-conditioned pick and place in clutter, where a robot should grasp a target object in open clutter and move it to a specified place. Some approaches learn end-to-end policies with features from vision foundation models, requiring large datasets. Others combine foundation models in a zero-shot setting, suffering from cascading errors. In addition, they primarily leverage vision and language foundation models, focusing less on action priors. In this paper, we aim to develop an effective policy by integrating foundation priors from vision, language, and action. We propose A2, an action prior alignment method that aligns unconditioned action priors with 3D vision-language priors by learning one attention layer. The alignment formulation enables our policy to train with less data and preserve zero-shot generalization capabilities. We show that a shared policy for both pick and place actions enhances the performance for each task, and introduce a policy adaptation scheme to accommodate the multi-modal nature of actions. Extensive experiments in simulation and the real-world show that our policy achieves higher task success rates with fewer steps for both pick and place tasks in clutter, effectively generalizing to unseen objects and language instructions. Videos and codes are available at the project page.
@@ I Introduction
Figure 1: Illustration of comparison on (a) classic end-to-end policies, (b) modular systems, and (c) our action prior alignment method that integrates foundation priors of vision, language, and action for planning.
@@ II-A Target-oriented Pick and Place in Clutter
Robotic pick and place in clutter has been a topic of interest in manipulation for decades. Traditional approaches [26, 27, 28, 29] are in the context of task and motion planning (TAMP) under the assumption of known object models and states. These methods struggle in open real scenarios, where obtaining precise object models and states is challenging. More recent research studies target-oriented unknown object grasping in clutter by first clearing obstacles [30, 31, 32], or retrieving the target object through non-prehensile actions [33, 34, 35, 1]. [2, 4, 6] step forward to build unknown object pick and place systems in cluttered environments, which are promising for real applications. However, these works still require images to specify target objects. Instead, language instructions are more flexible in open-world applications. By cooperating with foundation models, policies are capable of dealing with open-vocabulary objects in scattered scenes [7, 36, 9, 37]. In this paper, we aim to develop a policy for open-vocabulary pick and place in clutter, with the target specified with language instructions.
@@ II-B Foundation Models for Language-conditioned Manipulation
Foundation models in the field of CV and NLP have demonstrated powerful performance [38, 39, 40, 41, 42], and have been explored to facilitate robotic manipulation in open-world applications. A common way to utilize foundation models is to directly ground their capabilities into robotic scenarios. A series of approaches [13, 20, 43] uses vision foundation models for object grounding from flexible language instructions. Among them, some works explore object-centric representations for better scene understanding [13, 44, 45, 46, 47]. Other methods build 3D scene representations capturing both semantic and geometric information [48, 49, 50, 51, 52, 53]. For example, several approaches distill 3D neural feature fields from 2D foundation models [54, 19], requiring dense camera views and time-consuming training for high-quality rendering. This hinders real-time interaction in real-world scenarios. And efforts to overcome these limitations include introducing 3D Gaussian Splatting [16, 14, 15] and using sparse-view 3D representations [18, 17]. There are also methods [55, 21, 56, 22, 57, 58] utilizing the reasoning capability of large language models to build systems for planning. However, the performance of these policies largely depends on the capability of foundation models, and suffers from cascaded errors across individual modules. Another line of works [7, 8, 10, 9, 11, 12, 52] integrates features from vision foundation models into end-to-end policies. Despite promising results, these works consume extensive demonstration data and take plenty of training steps for convergence. In addition, one has to face the generalization issue if the tested objects or scenes are significantly different from those in the training data.
Recently, researchers have tried to learn action foundation models from large-scale robot data. For instance, AnyGrasp [59] is a grasp foundation model capable of generating grasp actions for open scenes. More generally, efforts are made to develop large Vision-Language-Action models (VLA) for general tasks and even embodiments [60, 61, 62, 63, 64]. However, leveraging priors from these action foundation models is much less explored. Some methods simply deploy pre-trained grasp models to generate grasp actions after object grounding, essentially paying less attention to the action planning side [19, 65]. In this paper, our policy aims to integrate priors from vision, language, and action foundation models to improve task performance.
@@ III Overview
Unconditioned Action Priors based Policy. Given the RGB-D image(s) =$\mathcal{I}\!=\!\{I_{i}\}_{i=0,1,...,M}$ and the language instruction $\mathcal{L}$, we leverage foundation models to extract vision, language, and action priors. Consider an action foundation model that generates $L$ candidate actions from image(s) as action priors $\mathcal{A}_{L}(\mathcal{I})=\{a_{k}\}_{k=0,1,...,L}$, $L$ generally has a controllable upper limit. These priors, distilled from a wide range of unconditioned data, provide feasible action patterns for downstream tasks and largely narrow the action space. Upon these priors, we construct a probabilistic policy $\pi$.
@@
$r(a,\mathcal{I},\mathcal{L})=\begin{cases}1,&\text{pick or place successfully}\\ 0,&\text{otherwise}\end{cases}\$ (2)
@@
$\max\omega\left(a_{d}|\mathcal{I}_{d},\mathcal{L}_{d}\right)$ (3)
@@
As a system, our policy first receives the language instruction to grasp the target object, and predicts a sequence of grasp actions by closed-loop action alignment. If the target cannot be directly grasped, our policy will select grasps around high-similarity regions to remove the obstacles and successfully pick up the target. Once the target is grasped, our policy takes the language instruction of placement along with the grasped object for place prediction, finally placing the grasped object in the assigned location.
@@ IV Foundation Priors
3D Representations. Given RGB-D image(s) =$\mathcal{I}\!=\!\{I_{i}\}_{i=0,1,...,M}$ from one or more cameras with fixed viewpoints, we first extract a 3D point cloud $\mathbf{p}$ within the workspace using the camera parameters. For each point $p_{j}$ of $\mathbf{p}$, we project it back to $i$th camera viewpoint as the pixel $u_{j}^{i}$, and get its visual feature $f_{j}^{i}$ by interpolation. Following [18], we compute weights for each camera according to the visibility and distance of $p_{j}$ in the corresponding camera. Finally, we fuse features from all camera viewpoints using a weighted sum, denoted as $f_{j}$. More details can be accessed in Appendix.
@@ V Action Prior Alignment
Based on foundation priors vision, language, and action, we propose to conduct action prior alignment by learning one attention layer.
@@ V-C Policy Learning
Flexible Language Instructions. Our policy is able to deal with flexible language instructions without assigning concrete object labels. For pick tasks, it can handle language instructions like “Give me the {target}” or “Get something to {target}”, where {target} can be a concrete label (e.g. banana), a general label (e.g. fruit), or even a functional description (e.g. hold other things). For place tasks, the language instructions are similar, but with additional spatial relation words, such as “Move the object {relation} the {reference}”. Here {relation} specifies the spatial relationship respective to the {reference}. {reference} is analogous to {target}, while {relation} can be words indicating “on” or “around” relations relative to reference}. For instance, words like “on top of”, and “into” belong to “on” relation, and others such as “next to”, and “near” belong to “around” relation.
@@ VI Implementation Details
Simulation Environment. We collect demonstration data by model-based expert planners with a UR5 arm in PyBullet [69]. There are three statically mounted Intel RealSense L515 cameras ($M=3$) overlooking the tabletop as shown in Fig. 2: one positioned 45°$\degree$ downward from the front, one 50°$\degree$ downward from the anti-diagonal perspective and one 50°$\degree$ downward from the diagonal perspective, referred to as the front, left and right cameras respectively. Our object models are from GraspNet-1Billion [24].
@@ VII-A Experimental Setup
Test Settings. We first conduct test experiments in simulation with a series of test cases, which can be categorized into three folds: pick, place, and pick-n-place. Each category includes cases of arrangements with both seen and unseen objects during A2 training. For place, some cases of unseen objects pair with unseen relations. Example cases are visualized in Fig. 3. For pick, each case contains 15 objects to form adversarial clutter where the robot might need to grasp away other obstacles for target grasping. All pick policies are evaluated on $x\!\!=\!\!10$ arrangements of seen objects and $x\!=\!5$ of unseen objects. For place, each case contains 8 objects to preserve some free space surrounding the reference object for placement. For place policies, performances are tested on $x\!\!=\!\!20$ arrangements of seen objects and $x\!\!=\!\!10$ of unseen objects. Given that placement is a one-step task, we add variances to action candidates in each test run of each case to evaluate robustness. For pick-n-place, we test policies with $x\!\!=\!\!8$ seen objects cases and $x\!\!=\!\!4$ unseen objects cases. Note that in these cases, we divide the workspace into a pick workspace (left) and a place workspace (right). To determine the success of target grasping, we use environment feedback in simulation. In real world, target is regarded as grasped if CLIP similarity of language and grasped object crop (filtered by depth) exceeds a threshold.
Evaluation Metrics. We evaluate the methods with a series of test cases. Each test contains $y=15$ runs measured with 2 metrics:
@@ VII-B Baselines
[…]
@@ VII-C Comparison to Baselines
Place. We show the performances of place with seen and unseen objects in Table I, further demonstrating the advantages of our action prior alignment paradigm. The performance of VLP depends on the capability of CLIP, which frequently fails when facing similar visual information or text words (e.g. “close to” and “to”). A2-G-Place struggles to distinguish “in” and “around” relation, as it directly grounds the highest point that fits both requirements of reference and relation.
@@ VII-D Ablation Studies
RL SL Res Data Seen Unseen
@@
TABLE V: Scaling to More Data and Object Number
@@
Figure 8: Test cases in real world. Each case contains 21$\sim$22 objects that are mostly unseen during training. Target or reference objects are labeled with stars.
Different Policy Adaptations. We compare different ways of policy adaptations, including different learning paradigms (RL, SL), data amounts and architectures (with or without the residual block). For RL fine-tuning, we deploy Soft Actor-Critic [73, 74]. For different architectures, Res represents whether using the residual block. If so, only the network parameters of the residual block are updated during fine-tuning. Otherwise, all the network parameters are fine-tuned. Performance comparisons on place tasks are reported in Table III. It is shown that fine-tuning with a residual block significantly outperforms full fine-tuning. This is because the residual block effectively injects multi-modality without losing much information of the pre-trained policy. With the residual block, both RL and SL can achieve comparable performances, whereas RL consumes much more data and demonstrates poorer generalization. This might be because place is a one-step task, where the lack of sequential decision making limits the effectiveness of RL optimization.
@@
Scaling to More Data and Object Number. We double the training data and report results as A2-2#D in Table. V, verifying effective improvements in both pick and place tasks when scaling to more data. To demonstrate more object distractors and dense clutter, we double the number of objects for testing as A2-2#O in Table V. Results show that our policy outperforms most baselines (tested with original object number) even with double objects, validating effectiveness in more complex settings.
Case Studies. Fig. 6 shows several cases in pick and place tasks to illustrate the 3D representations, the action priors, and the alignment results of our policy. Given the language instruction, the similarity cloud is able to highlight the task-relevant regions. Our policy aligns action priors based on zero-shot visual representations, enabling task-relevant action scoring for action selection. In addition, we show a case in Fig. 7, where direct visual grounding fails, and sampled points lack explicit shape, yet our policy selects a correct grasp through alignment. This highlights that although we take similarity-sampled points, we evaluate actions guided by visual grounding rather than being determined by it.
@@ VII-E Real-world Experiments
Figure 9: Real-world platform, which involves a UR5 robot arm equipped with a ROBOTIQ-85 gripper, and an Intel RealSense L515 capturing RGB-D images.
@@
Experiment Setup. In this section, we evaluate our policy in real-world settings. Fig. 9 shows our real-world setup, which involves a UR5 robot arm equipped with a ROBOTIQ-85 gripper, and an Intel RealSense L515 capturing RGB-D images at a resolution of 1280$\times$720. Notably, the camera viewpoint in the real-world setup is unseen during training. The workspace is divided into pick and place workspaces, where the robot is supposed to grasp the target object within the pick workspace, and then place it within place workspace. Our test cases consist of 5 scenarios shown in Fig. 8. Each of them contains 21$\sim$22 objects that are mostly unseen during training. There are in total of 38 objects for real-world testing, including 10 seen objects and 28 unseen objects.

### v3

@@ Abstract
We study the task of language-conditioned pick and place in clutter, where a robot should grasp a target object in open clutter and move it to a specified place. Some approaches learn end-to-end policies with features from vision foundation models, requiring large datasets. Others combine foundation models in a zero-shot setting, suffering from cascading errors. In addition, they primarily leverage vision and language foundation models, focusing less on action priors. In this paper, we aim to develop an effective policy by integrating foundation priors from vision, language, and action. We propose A2, an action prior alignment method that aligns unconditioned action priors with 3D vision-language priors by learning one attention layer. The alignment formulation enables our policy to train with less data and preserve zero-shot generalization capabilities. We show that a shared policy for both pick and place actions enhances the performance for each task, and introduce a policy adaptation scheme to accommodate the multi-modal nature of actions. Extensive experiments in simulation and the real-world show that our policy achieves higher task success rates with fewer steps for both pick and place tasks in clutter, effectively generalizing to unseen objects and language instructions. Videos and codes are available at https://xukechun.github.io/papers/A2.
Note to Practitioners
This research is motivated by the challenge of generalizable policy learning of language-conditioned pick and place in clutter. Solving such a challenge could significantly improve the robot’s level of automation and intelligence in household and industrial pick and place tasks. Existing methods struggle with large data requirements, poor generalization to unseen scenarios, and cascading errors across individual components. To overcome these limitations, we propose to integrate priors from vision, language, and action foundation models by learning-based alignment. Our policy aligns action priors with 3D vision-language priors by learning one attention layer, requiring less data and preserving zero-shot generalization capabilities from foundation models. Experiments show that our method can improve both task success rate and generalization for pick and place tasks in simulation and the real world. In future work, we will incorporate more action foundation models to extend our approach of action prior alignment to a wider range of tasks, offering a promising direction for general manipulation.
@@ I Introduction
Figure 1: Compared to previous methods (a) classic end-to-end policies and (b) modular systems, our method integrates foundation priors from vision, language, and action through alignment by one attention layer, which enables more efficient policy learning and better task performance.
@@ II-A Target-oriented Pick and Place in Clutter
Robotic pick and place in clutter has been a topic of interest in manipulation for decades. Traditional approaches [26, 27, 28, 29] are in the context of task and motion planning (TAMP) under the assumption of known object models and states. These methods struggle in open real scenarios, where obtaining precise object models and states is challenging. More recent research studies target-oriented unknown object grasping in clutter by first clearing obstacles [30, 31, 32], or retrieving the target object through non-prehensile actions [33, 34, 35, 1]. [2, 4, 6][36, 37] step forward to build unknown object pick and place systems in cluttered environments, which are promising for real applications. However, these works still require images to specify target objects. Instead, language instructions are more flexible in open-world applications. By cooperating with foundation models, policies are capable of dealing with open-vocabulary objects in scattered scenes [7, 38, 9, 39]. In this paper, we aim to develop a policy for open-vocabulary pick and place in clutter, with the target specified with language instructions.
@@ II-B Foundation Models for Language-conditioned Manipulation
Foundation models in the field of CV and NLP have demonstrated powerful performance [40, 41, 42, 43], and have been explored to facilitate robotic manipulation in open-world applications. A common way to utilize foundation models is to directly ground their capabilities into robotic scenarios. A series of approaches [13, 20, 44] uses vision foundation models for object grounding from flexible language instructions. Among them, some works explore object-centric representations for better scene understanding [13, 45, 46, 47, 48]. Other methods build 3D scene representations capturing both semantic and geometric information [49, 50, 51, 52, 53, 54]. For example, several approaches distill 3D neural feature fields from 2D foundation models [55, 19], requiring dense camera views and time-consuming training for high-quality rendering. This hinders real-time interaction in real-world scenarios. And efforts to overcome these limitations include introducing 3D Gaussian Splatting [16, 14, 15] and using sparse-view 3D representations [18, 17]. There are also methods [56, 21, 57, 22, 58, 59] utilizing the reasoning capability of large language models to build systems for planning. However, the performance of these policies largely depends on the capability of foundation models, and suffers from cascaded errors across individual modules. Another line of works [7, 8, 10, 9, 11, 12, 53] integrates features from vision foundation models into end-to-end policies. Despite promising results, these works consume extensive demonstration data and take plenty of training steps for convergence. In addition, one has to face the generalization issue if the tested objects or scenes are significantly different from those in the training data.
Recently, researchers have tried to learn action foundation models from large-scale robot data. For instance, AnyGrasp [60] is a grasp foundation model capable of generating grasp actions for open scenes. More generally, efforts are made to develop large Vision-Language-Action models (VLA) for general tasks and even embodiments [61, 62, 63, 64, 65]. However, leveraging priors from these action foundation models is much less explored. Some methods deploy pre-trained grasp models to generate grasp actions after object grounding, essentially paying less attention to action planning [19, 66]. In this paper, our policy aims to integrate priors from vision, language, and action foundation models to improve task performance.
@@ III Overview
Unconditioned Action Priors based Policy. Given the RGB-D image(s) $\mathcal{I}\!=\!\{I_{i}\}_{i=0,1,...,M}$ and the language instruction $\mathcal{L}$, we leverage foundation models to extract vision, language, and action priors. Consider an action foundation model that generates $L$ candidate actions from image(s) as action priors $\mathcal{A}_{L}(\mathcal{I})=\{a_{k}\}_{k=0,1,...,L}$, $L$ generally has a controllable upper limit. These priors, distilled from a wide range of unconditioned data, provide feasible action patterns for downstream tasks and largely narrow the action space. Upon these priors, we construct a probabilistic policy $\pi$.
@@
$r(a,\mathcal{I},\mathcal{L})=\begin{cases}1,&\text{pick or place successfully}\\ 0,&\text{otherwise}\end{cases}\\$ (2)
@@
$\max_{a_{d}\in\mathcal{A}_{L}(\mathcal{I}_{d})}\omega\left(a_{d}|\mathcal{I}_{d},\mathcal{L}_{d}\right)$ (3)
@@
As a system, our policy first receives the language instruction to grasp the target object, and predicts a sequence of grasp actions by closed-loop action alignment. If the target object is not grasped, our policy will remove the grasped obstacles and proceed regrasping. Once the target is successfully grasped, our policy takes the language instruction of placement along with the grasped object for place prediction, finally placing the grasped object in the assigned location.
@@ IV Foundation Priors
3D Representations. Given RGB-D image(s) $\mathcal{I}\!=\!\{I_{i}\}_{i=0,1,...,M}$ from one or more cameras with fixed viewpoints, we first extract a 3D point cloud $\mathbf{p}$ within the workspace using the camera parameters. For each point $p_{j}$ of $\mathbf{p}$, we project it back to $i$th camera viewpoint as the pixel $u_{j}^{i}$, and get its visual feature $f_{j}^{i}$ by interpolation. Following [18], we compute weights for each camera according to the visibility and distance of $p_{j}$ in the corresponding camera. Finally, we fuse features from all camera viewpoints using a weighted sum, denoted as $f_{j}$. More details can be accessed in Appendix.
@@ V Action Prior Alignment
Based on foundation priors, we propose to conduct action prior alignment by learning one attention layer.
@@ V-C Policy Learning
Flexible Language Instructions. Our policy is able to deal with flexible language instructions without assigning concrete object labels. For pick tasks, it can handle language instructions like “Give me the {target}” or “Get something to {target}”, where {target} can be a concrete label (e.g. banana), a general category (e.g. fruit), or the attribute of color (e.g. red), shape (e.g. round), or even a functional description (e.g. hold other things). For place tasks, the language instructions are similar, but with additional spatial relation words, such as “Move the object {relation} the {reference}”. Here {relation} specifies the spatial relationship respective to the {reference}. {reference} is analogous to {target}, while {relation} can be words indicating “on” or “around” relations relative to reference}. For instance, words like “on top of”, and “into” belong to “on” relation, and others such as “next to”, and “near” belong to “around” relation.
@@ VI Implementation Details
[…]
@@ VII-A Experimental Setup
Test Settings. We first conduct test experiments in simulation with a series of test cases, which can be categorized into three folds: pick, place, and pick-n-place. Each category includes cases of arrangements with both seen and unseen objects during A2 training. Specifically, seen objects are those that appear in the training set, while unseen objects are novel instances that are not observed during training. For place, some cases of unseen objects pair with unseen relations. Example cases are visualized in Fig. 3. For pick, each case contains 15 objects to form adversarial clutter where the robot might need to grasp away other obstacles for target grasping. All pick policies are evaluated on x=10x\!\!=\!\!10 arrangements of seen objects and x=5x\!=\!5 of unseen objects. For place, each case contains 8 objects to preserve some free space surrounding the reference object for placement. For place policies, performances are tested on x=20x\!\!=\!\!20 arrangements of seen objects and x=10x\!\!=\!\!10 of unseen objects. Given that placement is a one-step task, we add variances to action candidates in each test run of each case to evaluate robustness. For pick-n-place, we test policies with x=8x\!\!=\!\!8 seen objects cases and x=4x\!\!=\!\!4 unseen objects cases. Note that in these cases, we divide the workspace into a pick workspace (left) and a place workspace (right). To determine the success of target grasping, we use environment feedback in simulation. In real world, target is regarded as grasped if CLIP similarity of language and grasped object crop (filtered by depth) exceeds a threshold.
Evaluation Metrics. We evaluate methods with a series of test cases. Each contains y=15y\!=\!15 runs, measured with 2 metrics:
@@ VII-B Baselines
3D Visual Grounding Pick and Place Policies. We implement variant methods that directly conduct visual grounding using our 3D visual representations, named A2-G-Pick and A2-G-Place for pick and place tasks respectively. These methods select the action nearest to the region with the highest average similarity of K-nearest neighbors (K=0.05MM). In addition, A2-G-Pick can combine with A2-G-Place as a 3D visual grounding pick-n-place policy, denoted as A2-G.
@@ VII-C Comparison to Baselines
Place. We show the performances of place with seen and unseen objects in Table I, further demonstrating the advantages of our action prior alignment paradigm. The performance of VLP depends heavily on the capability of CLIP, which frequently fails when facing similar visual information or text words. A2-G-Place struggles to distinguish “in” and “around” relation, as it directly grounds the highest point that fits both requirements of reference and relation.
@@ VII-D Ablation Studies
RL IL Res Data Seen Unseen
@@
TABLE V: Scaling to More Objects and Data
@@
Figure 8: Example failure modes, including heavy occlusion, visual ambiguity, and semantic ambiguity.
@@
Different Policy Adaptations. We compare different policy adaptation strategies, including learning paradigms (reinforcement learning (RL), imitation learning (IL)), data amounts, and network architectures (with or without residual blocks). For RL, we adopt Soft Actor-Critic [74, 75]. In the “Res” variant, only the residual block is updated. Otherwise, the entire network is fine-tuned. As shown in Table III, residual-only fine-tuning significantly outperforms full fine-tuning, as it preserves pretrained knowledge while enabling efficient adaptation. In contrast, full fine-tuning may harm generalization, likely due to catastrophic forgetting. With residual blocks, IL and RL perform similarly, but RL requires more data and generalizes worse. This is likely because the one-step nature of place tasks limits RL’s advantage in sequential learning.
Network Architecture. We compare our method with some variant methods to evaluate the architecture design. Testing results are shown in Table IV. Removing RoPE causes notable drops: $17.3\%$ for unseen pick tasks and over $20\%$ for place tasks, highlighting its importance for generalization. In addition, using sampled point features as values in cross-attention instead of image features degrades performance, showing the benefit of foundation model features for effectiveness and generalization. Also, adding task-specific embeddings (TE) to action features harms performance, likely by hindering shared representations between pick and place. Finally, directly feeding the language embedding (LE) into cross-attention instead of weighting visual features with similarities weakens CLIP priors and reduces success rates.
@@
Generalization to More Objects. To demonstrate the generalization to more object distractors and denser clutter, we double the number of objects for testing as A2-2#O in Table V. Notably, there is no retraining of policy. Results show that our policy outperforms most baselines (tested with original object number) even with double objects, validating effectiveness in more complex settings.
Scaling to More Data. We double the training data to test the scalability, and report results as A2-2#D in Table. V, which verifies effective improvements in both pick and place tasks when scaling to more data.
Case Studies. Fig. 6 shows several cases to illustrate the 3D representations, action priors, and alignment results of our policy. Given language instructions, the similarity cloud can highlight the task-relevant regions, and our policy aligns action priors based on these representations. Fig. 7 further shows a case where visual grounding alone fails, but our alignment still enables correct grasp selection through alignment. This indicates that while we take similarity-sampled points, we evaluate actions guided by visual grounding rather than being determined by it. Fig. 8 visualizes some typical failure modes, including heavy occlusion and visual ambiguity of target objects, as well as semantic ambiguity in language instruction, i.e ambiguous word “cylinder”.
@@ VII-E Real-world Experiments
Figure 10: Real-world platform, which involves a UR5 robot arm equipped with a ROBOTIQ-85 gripper, and an Intel RealSense L515 camera.
@@
[…]